337 What is the empirical formula of a substance made up of 0.80 g of carbon and 0.20 g of hydrogen? 3.38 What is the empirical formula of a compound made up of 0.571 g of carbon, 0.072 g of hydrogen, and 0.333 g of nitrogen?
Answer #1

3.37)

we have mass of each elements as:

C: 0.8 g

H: 0.2 g

Divide by molar mass to get number of moles of each:

C: 0.8/12.01 = 0.0666

H: 0.2/1.008 = 0.1984

Divide by smallest to get simplest whole number ratio:

C: 0.0666/0.0666 = 1

H: 0.1984/0.0666 = 3

So empirical formula is:CH3

Answer: CH3